# Communication-Efficient and Privacy-Adaptable Mechanism for Federated Learning

### v2

@@ VI-B FL Convergence
We report the FL convergence of CEPAM-Gaussian in terms of accuracy using MLP and CNN architectures and MNIST dataset.
@@ VI-B1 Gaussian Mechanism
Figure 1: Convergence profile of different FL schemes.
In Table I, we report the test accuracy of CEPAM-Gaussian for different architectures (MLP and CNN) using MNIST with their 95% confidence intervals. CEPAM-Gaussian demonstrates better performance, achieving an improvement of 0.6-2.0% and 0.4-1.0% in accuracy compared to other baselines for MLP and CNN, respectively, suggesting that CEPAM can be beneficial for various learning models. Furthermore, all cases of CEPAM-Gaussian have approximately the same accuracy. This is because the added noise per dimension is consistent, i.e., $\mathrm{Var}(f)/n=\sigma^{2}$. However, higher-dimensional LRSUQ has a better compression ratio than scalar counterparts [21, 22] due to its nature as a vector quantizer. In contrast, for baselines such as the Gaussian-mechanism-then-quantize approach (Gaussian+SDQ), the total noise comprises privacy noise and quantization error, leading to more distortion and consequently lower accuracy compared to LRSUQ. It is worth noting that adding a minor level of distortion as a regularizer during training deep models can potentially enhance the performance of the converged model [78]. This finding is corroborated in our experimental results, where CEPAM-Gaussian achieves slightly better accuracy than vanilla FL for both architectures.
@@ A-A Proof of Lemma 11
Since we are using LRSUQ for the encoding and decoding of the $j$-th sub-vector $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$, where $j\in\mathcal{N}$, Proposition 5 implies that, regardless of the statistical model of $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$, the quantization errors $\{\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ are iid (over $k$ and $j$) and follows the pdf $f$.
@@ A-B Proof of Theorem 12
Our proof relies on the following fact about the mixture of uniform distributions [33].
@@
In other words, given the noise vector $\mathbf{Z}$, which follows a zero-mean multivariate Gaussian distribution with a covariance matrix $\sigma^{2}\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, i.e., $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{n})$, $\mathbf{Z}$ can be expressed as a mixture of uniform distributions over $n$-dimensional balls, with the latent variable $\tilde{U}$ following a chi-squared distribution with $n+2$ degrees of freedom. By setting $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ for every $k$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\;|\;\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for $k$ and $j$, Lemma 11 and Lemma 14 imply that the resulting mechanism, i.e., Step 7 in Algorithm 3, is Gaussian.
We utilize the subsampling with replacement mechanism to model the $\tau$-iterations of local SGD at $k$-th client, [^11] According to Equation (3), each $k$-th client performs $\tau$ iterations of local SGD with uniformly sampled indices from the dataset $\mathcal{D}^{(k)}$ (i.e., $p_{j}=1/|\mathcal{D}^{(k)}|=1/n$, where $j\in\{1,\ldots,n\}$ and $n=|\mathcal{D}^{(k)}|$ in (12) or (14)), this process can be regarded as subsampling with replacement, which outputs a multiset sample of size $\tau$ following a multinomial distribution $\mathrm{Mult}(\tau;p_{1},\ldots,p_{n})$. instead of employing Poisson subsampling for mini-batching [39] within a single local iteration as in [21]. Furthermore, the resulting subsampled Gaussian mechanism is applied to sub-vectors of dimension $n$ rather than to individual components.

### v3

@@ VI-B FL Convergence
We report the FL convergence in terms of accuracy for CEPAM-Gaussian using MLP and CNN architectures and CEPAM-Laplace using CNN architecture on MNIST dataset.
@@ VI-B1 Gaussian Mechanism
Figure 1: Convergence profile of different FL schemes for CEPAM-Gaussian.
In Table I, we report the test accuracy of CEPAM-Gaussian for different architectures (MLP and CNN) using MNIST with their 95% confidence intervals. CEPAM-Gaussian demonstrates better performance, achieving an improvement of 0.6-2.0% and 0.4-1.0% in accuracy compared to other baselines for MLP and CNN, respectively, suggesting that CEPAM can be beneficial for various learning models. Furthermore, all cases of CEPAM-Gaussian have approximately the same accuracy. This is because the added noise per dimension is consistent, i.e., $\mathrm{Var}(f)/n=\sigma^{2}$. However, higher-dimensional LRSUQ has a better compression ratio than scalar counterparts [hasircioglu2023communication, yan2023layered] due to its nature as a vector quantizer. In contrast, for baselines such as the Gaussian-mechanism-then-quantize approach (Gaussian+SDQ), the total noise comprises privacy noise and quantization error, leading to more distortion and consequently lower accuracy compared to LRSUQ. It is worth noting that adding a minor level of distortion as a regularizer during training deep models can potentially enhance the performance of the converged model [An1996AddingNoise]. This finding is corroborated in our experimental results, where CEPAM-Gaussian achieves slightly better accuracy than vanilla FL for both architectures.
@@
VI-B2 Laplace Mechanism
We set $b=0.001$ and $\tilde{\epsilon}$=30000 for the base Laplace mechanism in CEPAM-Laplace with a variance of $2\times 0.001^{2}$, and the composite Laplace mechanism achieves ($\epsilon=29995$)-DP. We also set $b=0.001$ for both FL+Laplace and FL+Laplace+SDQ. Note that a higher privacy budget is required to ensure $\epsilon$-DP for CEPAM-Laplace under privacy amplification with subsampling analysis. However, we might achieve a lower privacy budget by relaxing $\delta$. The trade-off is left for the future study.
Figure 2 shows the validation accuracy of CEPAM-Laplace over communication rounds using the CNN models, respectively. Similar to CEPAM-Gaussian, we observe that CEPAM-Laplace consistently outperforms other baselines.
Figure 2: Convergence profiles of different FL schemes for CEPAM-Laplace.
In Table II, we also report the performance of test accuracy of CEPAM-Laplace. We observe that CEPAM-Laplace achieves an improvement of about 1.5-1.6% in accuracy compared to other baselines for CNN, respectively, which suggests that CEPAM-Laplace can be beneficial for many learning models.
Table II: Test Accuracy (Laplace) for MNIST
Baselines CNN (%)
FL $92.92\pm 0.61$
FL+SDQ $92.93\pm 0.62$
FL+Laplace $92.95\pm 0.58$
FL+Laplace+SDQ $92.99\pm 0.57$
CEPAM-Laplace $94.53\pm 0.38$
@@ A-A Proof of Lemma 11
Since we are using LRSUQ for the encoding and decoding of the $j$-th sub-vector $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$ where $j\in\mathcal{N}$, Proposition 5 implies that, regardless of the statistical models of $\tilde{\mathbf{X}}_{t+\tau,j}^{k}$ where $j\in\mathcal{N}$, the quantization errors $\{\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\}_{j\in\mathcal{N}}$ are iid (over $k$ and $j$) and follows the pdf $f$.
@@ A-B Proof of Theorem 12
Our proof relies on the following fact about the mixture of uniform distributions [WalkerUniformP1999].
@@
In other words, given the noise vector $\mathbf{Z}$, which follows a zero-mean multivariate Gaussian distribution with a covariance matrix $\sigma^{2}\mathbf{I}_{n}$, where $\mathbf{I}_{n}$ is the $n\times n$ identity matrix, i.e., $\mathbf{Z}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I}_{n})$, $\mathbf{Z}$ can be expressed as a mixture of uniform distributions over $n$-dimensional balls, with the latent variable $\tilde{U}$ following a chi-squared distribution with $n+2$ degrees of freedom. By setting $\tilde{U}_{t+\tau,j}^{k}\sim g=\chi_{n+2}^{2}$ for every $k$ and $\tilde{\mathbf{Z}}_{t+\tau,j}^{k}\;|\;\{\tilde{U}_{t+\tau,j}^{k}=u\}\sim\mathrm{Unif}(\sigma\sqrt{u}B^{n})$ for $k$ and $j$, Lemma 11 and Lemma 14 imply that the resulting mechanism, i.e., Step 6 in Algorithm 3, is Gaussian.
We utilize the subsampling with replacement mechanism to model the $\tau$-iterations of local SGD at $k$-th client, [^11] According to Equation (3), each $k$-th client performs $\tau$ iterations of local SGD with uniformly sampled indices from the dataset $\mathcal{D}^{(k)}$ (i.e., $p_{j}=1/|\mathcal{D}^{(k)}|=1/n$, where $j\in\{1,\ldots,n\}$ and $n=|\mathcal{D}^{(k)}|$ in (12) or (14)), this process can be regarded as subsampling with replacement, which outputs a multiset sample of size $\tau$ following a multinomial distribution $\mathrm{Mult}(\tau;p_{1},\ldots,p_{n})$. instead of employing Poisson subsampling for mini-batching [Balle2018PrivAmpl] within a single local iteration as in [hasircioglu2023communication]. Furthermore, the resulting subsampled Gaussian mechanism is applied to sub-vectors of dimension $n$ rather than to individual components.